# When Should a Model Change Its Mind? An Energy-Based Theory and Regularizer for Concept Drift in Electrocardiogram (ECG) Signals

### v1

@@ Abstract
Models operating on dynamic physiologic signals must distinguish benign, label-preserving variability from true concept change. Existing concept-drift frameworks are largely distributional and provide no principled guidance on how much a model’s internal representation may move when the underlying signal undergoes physiologically plausible fluctuations in energy. As a result, deep models often misinterpret harmless changes in amplitude, rate, or morphology as concept drift, yielding unstable predictions—particularly in multimodal fusion settings. This study introduces Physiologic Energy Conservation Theory (PECT), an energy-based framework for concept stability in dynamic signals. PECT posits that under virtual drift, normalized latent displacement should scale proportionally with normalized signal-energy change, while persistent violations of this proportionality indicate real concept drift. We operationalize this principle through Energy-Constrained Representation Learning (ECRL), a lightweight regularizer that penalizes energy-inconsistent latent movement without modifying encoder architectures or adding inference-time cost. Although PECT is formulated for dynamic signals in general, we instantiate and evaluate it on multimodal ECG across seven unimodal and hybrid models. Experiments show that in the strongest trimodal hybrid (1D+2D+Transformer), clean accuracy is largely preserved ($96.0\%\rightarrow 94.1\%$), while perturbed accuracy improves substantially ($72.6\%\rightarrow 85.5\%$) and fused representation drift decreases by over $45\%$. Similar trends are observed across all architectures, providing empirical evidence that PECT functions as an energy–drift law governing concept stability in continuous physiologic signals.
@@ I Introduction
Most existing approaches to nonstationarity treat this problem through the lens of concept drift or, equivalently in biomedical settings, distribution/domain shift. Classical drift frameworks characterize change in terms of evolving distributions $P(X)$, $P(Y)$, or $P(Y\mid X)$ [6], and have been successfully applied in streaming and adaptive-learning scenarios. However, such formulations are largely agnostic to the physical structure of physiologic signals. In particular, they do not specify how much a learned representation is permitted to change when the underlying signal undergoes physiologically plausible, label-preserving variability. As a result, existing methods lack an explicit, physiologically grounded criterion for distinguishing virtual drift—benign variability that should preserve diagnostic meaning—from real concept drift, in which representations should legitimately traverse decision boundaries [17].
@@
This work addresses this gap by introducing Physiologic Energy Conservation Theory (PECT), an energy-based perspective on concept drift in dynamic physiologic signals. PECT posits that under label-preserving transformations, the magnitude of latent representation drift should scale proportionally with the corresponding change in signal energy. When representation displacement exceeds what physiologic energy variation can explain, the system transitions from virtual drift into real concept change. PECT thus provides a physically grounded criterion for determining when a model should change its mind, linking signal dynamics, representation geometry, and decision stability within a unified theoretical framework.
@@ I-A Contributions
    Decision stability via an energy-based drift rule. We formalize a principled decision-stability criterion that determines when a model should change its prediction based on violations of energy–drift proportionality, providing a deployable interpretation of concept drift beyond distributional change.
@@ II-D Physics-Informed and Mechanistic Learning
Physics-informed neural networks (PINNs) and related mechanistic learning approaches impose known physical constraints—often in the form of partial differential equations—on model outputs [14]. While philosophically aligned with PECT in their use of physical principles, these methods constrain outputs rather than representation dynamics. To our knowledge, no physics-informed framework uses measured physical or physiologic energy to regulate latent drift or to distinguish virtual from real concept drift in representation space.
@@ II-E Drift-Aware Representation Learning
Deep learning has achieved strong performance in ECG interpretation [7, 16]. Prior work on multimodal ECG fusion has examined domain complementarity, redundancy, and latent geometry [13, 12]. In particular, Complementary Feature Domain (CFD) theory provides an information-theoretic foundation for understanding how time, frequency, and time–frequency representations interact in multimodal ECG learning, while Physiologic Invariance Theory (PIT) characterizes robustness under label-preserving physiologic variability. PECT builds upon these foundations by introducing a distinct physical principle—energy-constrained drift—and extending it to the interpretation of concept drift in dynamic physiologic signals.
@@ II-G Comparison with Existing Frameworks
Table LABEL:tab:pect_theory_comparison summarizes how PECT and its operational formulation (ECRL) relate to existing theories and frameworks. In contrast to prior approaches, PECT uniquely links measured physiologic energy to latent representation drift and uses energy–drift proportionality as a criterion for distinguishing virtual drift from true concept change.
@@ III Physiologic Energy Conservation Theory
This section formalizes Physiologic Energy Conservation Theory (PECT), an energy-based framework for interpreting concept drift in dynamic physiologic signals. Rather than characterizing drift solely through prediction instability or distributional change, PECT describes how learned representations should respond to label-preserving physiologic energy fluctuations across individual modalities and their fused embedding. In concept-drift terms, PECT defines the expected behavior under virtual drift and establishes a boundary beyond which representation changes are more appropriately interpreted as true concept change.
@@ III-A Physiologic Energy in Cardiac Signals
which reflects aggregate electromechanical activity captured by the ECG waveform. Benign physiologic variations—such as mild amplitude scaling, heart-rate modulation, or temporal stretching—modulate $E(x)$ without altering diagnostic labels. In concept-drift terminology, such transformations correspond to virtual drift.
@@ III-B Energy Response of Multimodal Encoders
PECT posits that each encoder exhibits an intrinsic energy sensitivity that characterizes how latent features respond to physiologic energy variation:
$\delta z_{m}=g_{m}(\Delta E_{\mathrm{phys}}),$ (4)
@@ III-C Energy Consistency Principle and Concept Drift
PECT formalizes this requirement through the energy-consistency relation
$\|f_{m}(\tilde{x}_{m})-f_{m}(x_{m})\|_{2}\propto|\Delta E_{\mathrm{phys}}|.$ (5)
That is, latent displacement should scale proportionally with physiologic energy variation. When this proportionality is violated, identical energy shifts may produce disproportionately large latent movements, destabilizing fusion and degrading robustness.
This proportionality yields a natural concept-drift criterion: as long as (5) holds with a bounded proportionality constant, the observed drift is consistent with virtual drift. When latent displacement exceeds what physiologic energy variation can explain, the system enters a regime better interpreted as true concept change.
To quantify this relationship, PECT defines the modality-wise energy-coupling coefficient
$\mathcal{C}_{m}=\mathbb{E}_{x}\left[\frac{\|f_{m}(\tilde{x}_{m})-f_{m}(x_{m})\|_{2}}{|\Delta E_{\mathrm{phys}}|}\right],$ (6)
which is required to remain bounded for stable operation.
@@ III-D Energy Coupling in Multimodal Fusion
PECT requires that fused representation drift obey the same energy-consistency principle:
$z_{\mathrm{fused}}(\tilde{x})-z_{\mathrm{fused}}(x)=h(\Delta E_{\mathrm{phys}}),$ (8)
where $h(\cdot)$ captures the joint energy response across modalities.
@@
PECT Principle. Robust multimodal representations exhibit energy-proportional latent drift under label-preserving physiologic variability; violations of this proportionality indicate concept change. In relation to prior work, PECT extends Complementary Feature Domain (CFD) theory [13] and Physiologic Invariance Theory (PIT) by introducing an explicit physical constraint on how much multimodal representations are permitted to change under benign physiologic variability. Whereas CFD characterizes what information should be fused and PIT specifies which transformations should preserve predictions, PECT governs the magnitude of admissible representation drift induced by physiologic energy fluctuations.
@@ III-E Theoretical Justification of PECT
Assumption 1 (Energy-Lipschitz Encoders). Each modality-specific encoder satisfies
for all physiologic, label-preserving perturbations.
@@
Theorem 2 (Energy-Constrained Robustness Under Margins). If the fused representation drift satisfies
@@
then the predicted class remains invariant under physiologic perturbation for a margin-based classifier. In concept-drift terms, virtual drift does not alter the predicted label as long as energy-consistent representation displacement remains within the decision margin.
@@ IV-A Model Architecture Overview
To situate the proposed theory and training strategy within a concrete multimodal pipeline, Fig. 1 summarizes the complete architecture used throughout this study. The diagram illustrates how clean and physiologically perturbed ECG signals propagate through the temporal, time–frequency, and frequency-domain encoders, how their embeddings are fused into a joint latent representation, and where drift and physiologic energy responses are measured. This provides a structural link between PECT, the ECRL objective, and the role of each encoder in enforcing energy-proportional representation geometry.
@@ IV-B Dataset
We use a three-modality ECG dataset containing synchronized temporal, time–frequency, and frequency–domain representations derived from the same recordings. The dataset includes four diagnostic classes with clinically validated ground-truth labels. Raw signals are sampled at 500–1000 Hz and segmented into fixed-length windows spanning one to two cardiac cycles.
The dataset is split into training, validation, and held-out test sets using a patient-level partition to avoid sample leakage. After preprocessing (detailed below), the final dataset contains:
@@
  • multimodal alignment across all modalities.
@@ IV-I Implementation of PECT via ECRL
The theoretical framework developed in Section III is implemented through the Energy-Constrained Representation Learning (ECRL) procedure summarized in Algorithm 1. The algorithm integrates clean and physiologically perturbed samples, computes their energy differences, evaluates unimodal and fused representation shifts, and applies the BIT and PECT penalties to enforce energy-proportional drift. This provides a practical mechanism for translating the PECT hypothesis into a learnable objective that stabilizes multimodal representations under virtual drift.
@@
0: Minibatch $B=\{(x_{i},y_{i})\}$, physiologic perturbation operator $\mathcal{T}_{\text{phys}}$, encoders $f_{m}$, fusion module $F$, classifier $C$, weights $\lambda_{\text{BIT}}$, $\lambda_{\text{PECT}}$.
@@
2:  Generate physiologically perturbed input: $\tilde{x}_{i}=\mathcal{T}_{\text{phys}}(x_{i})$.
3:  Compute clean and perturbed physiologic energies: $E(x_{i})$ and $E(\tilde{x}_{i})$.
4:  Compute physiologic energy shift: $\Delta E_{\text{phys},i}=E(\tilde{x}_{i})-E(x_{i})$.
5:  Compute clean unimodal embeddings: $z^{(m)}_{\text{clean},i}=f_{m}(x_{i})$ for each modality $m$.
6:  Compute perturbed unimodal embeddings: $z^{(m)}_{\text{pert},i}=f_{m}(\tilde{x}_{i})$ for each modality $m$.
7:  Fuse clean embeddings: $z^{\text{clean}}_{\text{fused},i}=F(\{z^{(m)}_{\text{clean},i}\})$.
8:  Fuse perturbed embeddings: $z^{\text{pert}}_{\text{fused},i}=F(\{z^{(m)}_{\text{pert},i}\})$.
9:  Compute fused drift: $\delta z_{\text{fused},i}=z^{\text{pert}}_{\text{fused},i}-z^{\text{clean}}_{\text{fused},i}$.
10:  Compute energy–coupling coefficient: $\Gamma_{i}=\|\delta z_{\text{fused},i}\|_{2}\,/\,|\Delta E_{\text{phys},i}|$.
11:  Compute classification loss: $\mathcal{L}^{(i)}_{\text{cls}}=\ell\!\big(C(z^{\text{clean}}_{\text{fused},i}),y_{i}\big)$.
12:  Compute BIT loss: $\mathcal{L}^{(i)}_{\text{BIT}}=\|\delta z_{\text{fused},i}\|_{2}^{2}$.
13:  Compute PECT loss: $\mathcal{L}^{(i)}_{\text{PECT}}=(\Gamma_{i}-\mu_{\Gamma})^{2}$.
14:  Compute per-sample ECRL objective: $L^{(i)}=\mathcal{L}^{(i)}_{\text{cls}}+\lambda_{\text{BIT}}\mathcal{L}^{(i)}_{\text{BIT}}+\lambda_{\text{PECT}}\mathcal{L}^{(i)}_{\text{PECT}}$.
15:end for
16: Aggregate minibatch loss: $\mathcal{L}_{\text{ECRL}}=\frac{1}{|B|}\sum_{i}L^{(i)}$.
17: Update fusion + classifier parameters via $\nabla_{\theta}\mathcal{L}_{\text{ECRL}}$ (e.g., Adam).

### v2

# arXiv:2602.22294v2[cs.LG] 27 Feb 2026

@@ Abstract
Models operating on dynamic physiologic signals must distinguish benign, label-preserving variability from true concept change. Existing concept-drift frameworks are largely distributional and provide no principled guidance on how much a model’s internal representation may move when the underlying signal undergoes physiologically plausible fluctuations in energy. As a result, deep models often misinterpret harmless changes in amplitude, rate, or morphology as concept drift, yielding unstable predictions, particularly in multimodal fusion settings. This study introduces Physiologic Energy Conservation Theory (PECT), an energy-based framework for concept stability in dynamic signals. PECT posits that under virtual drift, normalized latent displacement should scale approximately proportionally with normalized signal-energy change, while persistent violations of this relationship indicate real concept drift. We operationalize this principle through Energy-Constrained Representation Learning (ECRL), a lightweight regularizer that penalizes energy-inconsistent latent movement without modifying encoder architectures or adding inference-time cost. Although PECT is formulated for dynamic signals in general, we instantiate and evaluate it on multimodal ECG across seven unimodal and hybrid models. Experiments show that in the strongest trimodal hybrid (1D+2D+Transformer), clean accuracy is largely preserved ($96.0\%\rightarrow 94.1\%$), while perturbed accuracy improves substantially ($72.6\%\rightarrow 85.5\%$) and fused representation drift decreases by over $45\%$. Similar trends are observed across all architectures, providing empirical evidence that PECT functions as an energy–drift law governing concept stability in continuous physiologic signals.
@@ I Introduction
Most existing approaches to nonstationarity treat this problem through the lens of concept drift or, equivalently in biomedical settings, distribution/domain shift. Classical drift frameworks characterize change in terms of evolving distributions $P(X)$, $P(Y)$, or $P(Y\mid X)$ [6], and have been successfully applied in streaming and adaptive-learning scenarios. However, such formulations are largely agnostic to the physical structure of physiologic signals. In particular, they do not specify how much a learned representation is permitted to change when the underlying signal undergoes physiologically plausible, label-preserving variability. As a result, existing methods lack an explicit, physiologically grounded criterion for distinguishing virtual drift (benign variability that should preserve diagnostic meaning) from real concept drift, in which representations should legitimately traverse decision boundaries [19].
@@
This work addresses this gap by introducing Physiologic Energy Conservation Theory (PECT), an energy-based perspective on concept drift in dynamic physiologic signals. PECT posits that under label-preserving transformations, the magnitude of latent representation drift should scale approximately proportionally with the corresponding change in signal energy. When representation displacement exceeds what physiologic energy variation can explain, the system transitions from virtual drift into real concept change. PECT thus provides a physically grounded criterion for determining when a model should change its mind, linking signal dynamics, representation geometry, and decision stability within a unified theoretical framework.
@@ I-A Contributions
    Decision stability via an energy-based drift rule. We formalize a principled decision-stability criterion that determines when a model should change its prediction based on violations of energy–drift consistency, providing a deployable interpretation of concept drift beyond distributional change.
@@ II-D Physics-Informed and Mechanistic Learning
Physics-informed neural networks (PINNs) and related mechanistic learning approaches impose known physical constraints (often in the form of partial differential equations) on model outputs [16]. While philosophically aligned with PECT in their use of physical principles, these methods constrain outputs rather than representation dynamics. To our knowledge, no physics-informed framework uses measured physical or physiologic energy to regulate latent drift or to distinguish virtual from real concept drift in representation space.
@@ II-E Drift-Aware Representation Learning
Deep learning has achieved strong performance in ECG interpretation [7, 18]. Prior work on multimodal ECG fusion has examined domain complementarity, redundancy, and latent geometry [15, 13]. In particular, Complementary Feature Domain (CFD) theory provides an information-theoretic foundation for understanding how time, frequency, and time–frequency representations interact in multimodal ECG learning, while Physiologic Invariance Theory (PIT) characterizes robustness under label-preserving physiologic variability [14]. PECT builds upon these foundations by introducing a distinct physical principle—energy-constrained drift—and extending it to the interpretation of concept drift in dynamic physiologic signals.
@@ II-G Comparison with Existing Frameworks
Table I summarizes how PECT and its operational formulation (ECRL) relate to existing theories and frameworks. In contrast to prior approaches, PECT uniquely links measured physiologic energy to latent representation drift and uses energy–drift consistency as a criterion for distinguishing virtual drift from true concept change.
TABLE I: How PECT/ECRL differs from existing stability and drift frameworks. PECT uniquely ties measured physiologic signal energy to latent representation drift and uses energy–drift consistency to separate virtual drift from real concept change.
Framework family Primary object of change Stability notion Uses measured signal energy? What it cannot do (gap PECT fills)
Concept-drift / distributional drift [6, 12] Input/label distributions ($P(X)$, $P(Y)$, $P(Y|X)$) Detect/track distribution change; trigger adaptation No No principled bound on how much latent space may move under label-preserving physiologic variability.
Robustness / Lipschitz / margin bounds [2, 1, 20] Model mapping sensitivity (input $\rightarrow$ features/logits) Small input change $\Rightarrow$ bounded output/feature change No Perturbations are abstract (e.g., norm-bounded); does not relate drift to a physiologically interpretable quantity like $\Delta E_{\text{phys}}$.
Invariance / equivariance (group-based) [3, 8] Transformation structure (symmetries) Representation should be invariant/equivariant to certain transforms No Specifies which transforms should preserve outputs, but not the magnitude law governing admissible latent motion under physiologic variability.
Energy-Based Models (EBMs) [11, 4] Learned scalar energy function over configurations Low learned energy for plausible states No (learned energy) “Energy” is a learned potential, not physical/physiologic waveform energy; no energy–drift consistency constraint in representation space.
Physics-informed learning (PINNs, mechanistic constraints) [16] Outputs constrained by known physics (often PDEs) Physical consistency of outputs/trajectories Sometimes (via equations) Typically constrains outputs, not latent drift geometry; does not define a runtime criterion separating virtual drift from real drift via $\|\delta z\|$ and $|\Delta E|$.
Continual learning / representation-drift penalties [10, 21] Feature/parameter drift across tasks or time Reduce forgetting; stabilize features across domains/tasks No Targets task/domain shifts rather than benign within-class physiologic variability; no physically grounded admissible-drift rule.
PECT (this work) Latent drift as a function of physiologic energy Virtual drift exhibits energy-consistent latent motion; persistent violations indicate concept change Yes Provides a physically interpretable drift law and a deployable criterion for “when a model should change its mind.”
ECRL (this work) Train-time regularizer enforcing PECT Penalizes energy-inconsistent latent motion (architecture-agnostic) Yes Stabilizes unimodal and fused embeddings under physiologic perturbations without inference-time cost.
@@ III Physiologic Energy Conservation Theory
This section formalizes Physiologic Energy Conservation Theory (PECT), an energy-based framework for interpreting concept drift in dynamic physiologic signals. Rather than characterizing drift solely through prediction instability or distributional change, PECT describes how learned representations should respond to label-preserving physiologic energy fluctuations across individual modalities and their fused embedding. In concept-drift terms, PECT defines expected behavior under virtual drift and establishes a boundary beyond which representation changes are more appropriately interpreted as true concept change.
@@ III-A Physiologic Energy in Cardiac Signals
which reflects aggregate electromechanical activity captured by the ECG waveform. In a discrete implementation, energy can be approximated by $E(x)=\Delta t\sum_{n=1}^{N}x[n]^{2}$ (or $\sum_{n=1}^{N}x[n]^{2}$ after normalization). Benign physiologic variations—such as mild amplitude scaling, heart-rate modulation, or temporal stretching—modulate $E(x)$ without altering diagnostic labels. In concept-drift terminology, such transformations correspond to virtual drift.
@@ III-B Energy Response of Multimodal Encoders
For physiologic perturbations $\tilde{x}_{m}=\mathcal{T}_{\text{phys}}(x_{m})$, define latent drift
$\delta z_{m}=f_{m}(\tilde{x}_{m})-f_{m}(x_{m}).$ (4)
PECT characterizes modality-specific energy sensitivity through the magnitude of $\delta z_{m}$. Under virtual drift, PECT predicts that the expected latent displacement conditioned on energy change is bounded and approximately monotone:
$\mathbb{E}\!\left[\|\delta z_{m}\|_{2}\,\middle|\,|\Delta E_{\mathrm{phys}}|\right]\approx g_{m}\!\left(|\Delta E_{\mathrm{phys}}|\right),$ (5)
@@ III-C Energy Consistency Principle and Concept Drift
PECT formalizes this requirement through an energy-Lipschitz condition:
for all physiologic, label-preserving perturbations in $\mathcal{T}_{\text{phys}}$. When this relationship is persistently violated, identical energy shifts may produce disproportionately large latent movements, destabilizing fusion and degrading robustness.
This yields a concept-drift criterion: as long as (6) holds with bounded constant $L_{m}$, the observed drift is consistent with virtual drift. When latent displacement exceeds what physiologic energy variation can explain, the system enters a regime better interpreted as true concept change.
To quantify coupling, PECT defines the modality-wise energy-coupling coefficient
$\mathcal{C}_{m}=\mathbb{E}_{x}\left[\frac{\|f_{m}(\tilde{x}_{m})-f_{m}(x_{m})\|_{2}}{|\Delta E_{\mathrm{phys}}|+\epsilon}\right],$ (7)
where $\epsilon>0$ prevents degeneracy when $|\Delta E_{\mathrm{phys}}|\approx 0$.
@@ III-D Energy Coupling in Multimodal Fusion
Define fused drift $\delta z_{\mathrm{fused}}=z_{\mathrm{fused}}(\tilde{x})-z_{\mathrm{fused}}(x)$. PECT requires that fused representation drift obey an energy-consistency principle:
$\|\delta z_{\mathrm{fused}}\|_{2}\leq\kappa\,|\Delta E_{\mathrm{phys}}|.$ (9)
@@
PECT Principle. Robust multimodal representations exhibit energy-consistent latent motion under label-preserving physiologic variability; persistent violations indicate concept change. In relation to prior work, PECT extends Complementary Feature Domain (CFD) theory [15] and Physiologic Invariance Theory (PIT) [14] by introducing an explicit physical constraint on how much multimodal representations are permitted to change under benign physiologic variability.
@@ III-E Theoretical Justification of PECT
Assumption 1 (Energy-Lipschitz Encoders). Each modality-specific encoder satisfies (6) for all physiologic, label-preserving perturbations.
@@
Theorem 2 (Energy-Constrained Robustness Under Margins). Let $C$ be a margin-based classifier in fused space with (fused-space) margin $\gamma(x)>0$ at $z_{\mathrm{fused}}(x)$, where $\gamma(x)$ denotes the minimum distance from $z_{\mathrm{fused}}(x)$ to the decision boundary in the $\|\cdot\|_{2}$ metric. If the fused representation drift satisfies
@@
then the predicted class remains invariant under physiologic perturbation.
@@ IV-A Model Architecture Overview
To situate the proposed theory and training strategy within a concrete multimodal pipeline, Fig. 1 summarizes the complete architecture used throughout this study. The diagram illustrates how clean and physiologically perturbed ECG signals propagate through the temporal, time–frequency, and frequency-domain encoders, how their embeddings are fused into a joint latent representation, and where drift and physiologic energy responses are measured. This provides a structural link between PECT, the ECRL objective, and the role of each encoder in enforcing energy-consistent representation geometry.
@@ IV-B Dataset
We used the ECG dataset of cardiac patients made publicly available on Mendeley Data by Khan and Hussain [9]. From each recording, we generated synchronized temporal (1D), frequency-domain (1D), and time–frequency (2D) representations, ensuring that all modalities correspond to the same underlying cardiac cycle. The dataset includes four diagnostic classes with clinically validated ground-truth labels. The raw signals are sampled at 500–1000 Hz and segmented into fixed-length windows spanning one to two cardiac cycles.
The dataset is divided into training, validation, and held-out test sets using a patient-level partition to prevent sample leakage across splits. After preprocessing (described in Section IV-C), the final dataset contains
@@
  • fully aligned samples across all three modalities.
@@ IV-I Implementation of PECT via ECRL
The theoretical framework developed in Section III is implemented through the Energy-Constrained Representation Learning (ECRL) procedure summarized in Algorithm 1. The algorithm integrates clean and physiologically perturbed samples, computes their energy differences, evaluates unimodal and fused representation shifts, and applies the BIT and PECT penalties to enforce energy-consistent drift.
@@
0: Minibatch $B=\{(x_{i},y_{i})\}$, physiologic perturbation operator $\mathcal{T}_{\text{phys}}$, encoders $f_{m}$, fusion module $F$, classifier $C$, weights $\lambda_{\text{BIT}}$, $\lambda_{\text{PECT}}$, stability constant $\epsilon>0$.
@@
2:  Generate perturbed input: $\tilde{x}_{i}=\mathcal{T}_{\text{phys}}(x_{i})$.
3:  Compute energies $E(x_{i})$ and $E(\tilde{x}_{i})$ and energy shift $\Delta E_{\text{phys},i}=E(\tilde{x}_{i})-E(x_{i})$.
4:  Compute clean embeddings $z^{(m)}_{\text{clean},i}=f_{m}(x_{i})$ and perturbed embeddings $z^{(m)}_{\text{pert},i}=f_{m}(\tilde{x}_{i})$.
5:  Fuse clean embeddings: $z^{\text{clean}}_{\text{fused},i}=F(\{z^{(m)}_{\text{clean},i}\})$.
6:  Fuse perturbed embeddings: $z^{\text{pert}}_{\text{fused},i}=F(\{z^{(m)}_{\text{pert},i}\})$.
7:  Compute fused drift: $\delta z_{\text{fused},i}=z^{\text{pert}}_{\text{fused},i}-z^{\text{clean}}_{\text{fused},i}$.
8:  Compute energy–coupling coefficient: $\Gamma_{i}=\|\delta z_{\text{fused},i}\|_{2}\,/\,(|\Delta E_{\text{phys},i}|+\epsilon)$.
9:end for
10: Compute batch target $\mu_{\Gamma}=\frac{1}{|B|}\sum_{i\in B}\Gamma_{i}$.
11:for each $(x_{i},y_{i})$ in $B$ do
12:  Classification loss: $\mathcal{L}^{(i)}_{\text{cls}}=\ell\!\big(C(z^{\text{clean}}_{\text{fused},i}),y_{i}\big)$.
13:  BIT loss: $\mathcal{L}^{(i)}_{\text{BIT}}=\|\delta z_{\text{fused},i}\|_{2}^{2}$.
14:  PECT loss: $\mathcal{L}^{(i)}_{\text{PECT}}=(\Gamma_{i}-\mu_{\Gamma})^{2}$.
15:  Per-sample objective: $L^{(i)}=\mathcal{L}^{(i)}_{\text{cls}}+\lambda_{\text{BIT}}\mathcal{L}^{(i)}_{\text{BIT}}+\lambda_{\text{PECT}}\mathcal{L}^{(i)}_{\text{PECT}}$.
16:end for
17: Aggregate minibatch loss: $\mathcal{L}_{\text{ECRL}}=\frac{1}{|B|}\sum_{i\in B}L^{(i)}$.
18: Update parameters via $\nabla_{\theta}\mathcal{L}_{\text{ECRL}}$ (e.g., Adam).